Lucky Boys Auto's

Lucky Boys Autos is considering the way it approaches the market. First, it can sell cars by posting a “always low (fixed) price. Customers who are willing to buy simply complete paperwork and the sale is complete.

Alternatively, Boys can hire commissioned sales reps to probe customers willingness and ability to pay. (The reps are trained to ask customers where they live and work, and often directly ask them what they intend to spend!).

Under the uniform pricing strategy, the firm's MC is $2 (thousand) per unit (which is also the AVC). The latter strategy increases MC and AVC by $1 thousand per car (to $3) but face FC of $6K per month regardless of the number of cars sold. Here, each customer pays their full reservation price.

With a demand curve of P = 10 - Q, which strategy should Lucky Boys use and why?
Answer #1

A monopolist always maximizes its profit at MR=MC .

Demand Curve : P=10-Q Profit-maximizing condition : MR = Me MR = d(TR) = d (P. Q) d (10Q - Q = 10-20 aq de de hen me = a 10-2etive When Ave=3 , there is a fixed cost of 6 wore Of Q Hence, TC = TVC+TFC = 3Q+6 T = 100-Q _ 30-6 = 70-07-6 - 7 (7/2)-(7/2)